OpenStudy (anonymous):

Neil bought a refrigerator from a rental center for $1,519. He makes six monthly payments of $287.35 with his credit card. The rental center charges $3.35 for every purchase made with a credit card. He also had to pay late fees in the amount of $37 three different times. What is the total cost of the refrigerator? $1,855.20 $1,744.20 $1,835.10 $1,724.10

OpenStudy (jack1):

6 payments of 287.35 + 6 fees @ 3.35 + 3 late fees of 37 = ???

OpenStudy (jack1):

@nadiiraa_k ?

OpenStudy (anonymous):

I got the answer of 1855.20 But i'm unaware if its correct.

OpenStudy (anonymous):

My boyfriend fot 1835.10

OpenStudy (jack1):

hang on, i just have to do the math, wont be a sec

OpenStudy (jack1):

6 payments of 287.35 = 1724.1 + 6 fees @ 3.35 = 20.1 + 3 late fees of 37 = 111 sum = $1855.20

OpenStudy (jack1):

no, you're correct, he's wrong

OpenStudy (jack1):

wait

OpenStudy (jack1):

"The rental center charges $3.35 for every purchase made with a credit card" does this mean for every payment made on his card? or just for the original purchase?

OpenStudy (anonymous):

I'm not sure, it doesn't specify. But i'm assuming that because he uses the card to make the payment it would count right?

OpenStudy (jack1):

yeah, i'm just not sure now if i means we count it once or 6 times tho...??? hmmm, sorry, I think our answers correct still, but question whoever set the q's for u maybe?

OpenStudy (anonymous):

Its a exam for my online class, I'd ask my teacher but I doubt he'd answer.. :(

OpenStudy (anonymous):

I'll just go with the answer we got!

OpenStudy (jack1):

cools, best of luck